OpenStudy (anonymous):

in a chess match, a win counts 1 point, a draw counts 1/2 point, and a loss counts 0 points. After 15 games, the winner was 4 points ahead of the loser. How many points did the loser have?

OpenStudy (anonymous):

5.5

OpenStudy (anonymous):

how u do that?

OpenStudy (anonymous):

suppose the winner won the first four matches and all the rest were drawn, then 4 matches were won and 11 were drawn. so the loser has 11 times 1/2 points = 5.5 points

OpenStudy (anonymous):

actually there is a better explanation. let me know if you want it

OpenStudy (anonymous):

want it!

OpenStudy (anonymous):

okay. the winner is 4 points ahead of the loser. so the winner won 4 more games than the loser. now, there are two possibilities which explain this scenario. 1) the winner won 4 games and the rest were all drawn. 2) the winner won 9 games, the loser won 5 games and there was one draw. in each case the loser has 5.5 points.

OpenStudy (anonymous):

oh... geddit. tnx

OpenStudy (anonymous):

hmm. winner 4 points ahead of loser. loser has x, winner has x+4 \[x+x+4=15\] \[2x+4=15\] \[2x=11\] \[x=5.5\] loser has 5.5, winner has 9.5 no matter what happened in each game.